# Realization of the Tellegen Effect in Resonant Optical Metasurfaces

### v1

@@ Abstract
We measure the magneto-optical Kerr rotation angle ($\theta$) and Kerr ellipticity ($\varepsilon$) using a magneto-optical setup illustrated in Fig. 3(c) (see more details in Methods and Supplementary Section 6). The samples were positioned between the poles of an electromagnet generating a magnetic field of $B=\pm 1.3$ T, perpendicular to the measured surface. This field alternately reversed the spontaneous magnetization of the nanocones along their axes. The obtained hysteresis curves can be found in Supplementary Section 7. The incident light was directed normally onto the metasurface, passing through a hole in the electromagnet. In the measurements of the Tellegen effect, magneto-optical data are acquired under an external magnetic field of $B=1.3$ T. This ensures that the meta-atoms in all three metasurfaces remain in the same uniformly saturated magnetization state, making the metasurfaces magneto-optically equivalent and enabling the use of our extraction technique. Measuring the Tellegen effect in the absence of an external field is also possible but would require more costly and far less scalable nanofabrication that results in identical meta-atoms diameters across the metasurface.
@@
Figures 3(d)–(f) present the experimental Kerr rotation and ellipticity for the three metasurfaces ($\theta_{1,2,3}$ and $\epsilon_{1,2,3}$) as functions of the free-space wavelength. For comparison, we also include the simulated Kerr parameters for periodic metasurfaces. It is important to note that these full-wave simulations cannot fully capture the complexity of the experimental metasurfaces, featuring meta-atoms with significant variations in their overall sizes (although with the same proportions) and quasi-random distances between neighboring meta-atoms. In the simulations, we assume a periodic arrangement of meta-atoms with an effective period of $a=330$ nm, using averaged geometric parameters of the meta-atoms depicted in the inset of Fig. 3b).
@@
Despite the inherent differences between the quasi-random experimental metasurfaces and their periodic simulated counterparts, we observe both qualitative and satisfactory quantitative agreement between the measured and simulated Kerr spectra. All the spectra exhibit the characteristic Lorentzian dispersion. For the first metasurface with a spacer thickness of $h_{1}$, the simulated Kerr parameters are slightly larger in amplitude than the measured ones, although the spectral correspondence remains strong. For the second metasurface with thickness $h_{2}$, the experimental and simulated data show good agreement both spectrally and quantitatively. However, for the third metasurface with $h_{3}$, the simulated results are redshifted by approximately 150 nm relative to the experimental data.
@@
By applying our proposed extraction technique (summarized in Fig. 2), we calculate the polarizabilities of the experimental metasurfaces using both measured and simulated data from Figs. 3(d)–(f). The extracted polarizability spectra are plotted in Figs. 3(g)–(i). One can observe a good qualitative agreement between the measured and simulated data, both spectrally and in terms of amplitude. All three effects—gyroelectric, gyromagnetic, and Tellegen—exhibit comparable amplitudes within the wavelength range of interest. Comparing these results with the simulated polarizabilities of the idealized meta-atom in Fig. 1, we observe similar response, although the resonance frequencies of the effects have shifted. These quantitative differences stem from the substantial geometric variations between the idealized meta-atom in Fig. 1(a) and the realistic meta-atom in Fig. 3(b). Furthermore, in our extraction, we model the array of meta-atoms as a homogenized sheet of electric and magnetic currents with deeply subwavelength thickness. This introduces an additional source of inaccuracy, as the actual meta-atom height is on the order of $\lambda/6$. Nevertheless, both the simulated and measured results provide strong evidence for the resonant nature of the diagonal Tellegen effect in the proposed optical metasurface. The magnitudes of the Tellegen polarizability $|\alpha_{\rm te}|$, extracted from both the measured data and the simulations for the idealized meta-atoms in Fig. 1 at 830 nm, reaches approximately $6\times 10^{-33}$ ${\rm m}^{2}\cdot{\rm s}$. These results demonstrate that the observed amplitude of the Tellegen effect is comparable to that of the resonant gyroelectric effect present in the metasurface. This implies that the Tellegen effect is sufficiently strong to be relevant for magneto-optical applications. In the following section, we calculate and discuss the magnitude of the observed value and determine the corresponding bulk magnetoelectric parameter.
Discussion
@@
To evaluate the macroscopic effective-medium parameters of the purely Tellegen colloid shown in Fig. 4(d), we apply the Maxwell-Garnett mixing rule [15, Sec. 7.2.2]. Here, we assume the host medium to be water, however, other dielectric media would yield comparable results, differing mainly in the effective permittivity of the colloid. The volumetric concentration of the meta-atoms is set to $N=3.47\times 10^{20}\leavevmode\nobreak\ {\rm m}^{-3}$, corresponding to an experimentally feasible volume fraction of the cones $N_{V}=0.1$. Figure 4(e) presents the effective bulk calculated Tellegen parameter, $\chi_{\rm eff}$, exhibiting a Lorentzian resonance spectral shape. At resonance near 850 nm, $\chi_{\rm eff}$ reaches approximately $10^{-3}$, which is at least 100 times larger than values observed in natural materials such as chromium oxide [47]. When using the experimentally calculated Tellegen polarizability from Fig. 3(i) in our theoretical calculations, we obtain a slightly smaller value of the Tellegen bulk parameter, $\chi_{\rm eff}=8\times 10^{-4}$, at the resonance. Owing to its isotropic response, the Tellegen colloid exhibits identical magneto-optical behavior regardless of the illumination direction. Figure 4(f) plots these “isotropic” Kerr parameters for the colloid, showing qualitative similarity to those of the membrane in Fig. 4(c), as both systems exhibit a purely Tellegen response.
@@
In conclusion, we have demonstrated the first experimental realization of the optical Tellegen effect in artificial metamaterial system, with the strength of around 100 times greater than that of natural materials. The Tellegen metasurface was produced using conventional materials and a well-established, highly-scalable nanofabrication, ensuring practicality for diverse applications. Since the presented optical metasurface also exhibits gyroelectric and gyromagnetic effects, we introduced an approach for independently extracting each of the three nonreciprocal effects. This technique is expected to be instrumental in the magneto-optical characterization of other nonreciprocal optical metasurfaces, particularly where symmetry permits multiple nonreciprocal effects. Experimental measurements, including those of magneto-optical Kerr parameters and nonreciprocal polarizabilities, demonstrate strong agreement with full-wave simulations, confirming the accuracy and reliability of the proposed method.
@@
The inherent scalability of the chosen bottom-up nanofabrication approach facilitates the straightforward extension of the proposed nonreciprocal optical Tellegen metasurface to other metamaterial systems, including bulk Tellegen metamaterials and colloids. By tailoring the geometry and composition of meta-atoms, the resonant wavelength can be precisely tuned across the optical spectrum. Further, opening-up the ability to modulate the Tellegen metamaterial in space and/or time has profound implications for generating and analyzing effective axion fields— which are inherently nonuniform [54, 55, 56]— and other related phenomena [2, 3, 4, 57, 58].
Materials and methods
Simulation
@@
Magneto-optical characterization
The magneto-optical Kerr effect (MOKE) of the samples was characterized using an NKT SuperK EXW-12 supercontinuum laser connected to an acousto-optical tunable filter (AOTF), providing a wavelength range of 500–1400 nm. The AOTF employed in our setup has two outputs: VIS/NIR (500–900 nm) and NIR2 (800–1400 nm). A focusing doublet, Glan-Thompson polarizer, and silver D-shaped mirror were used to direct the light into the magnet poles and focus it onto the samples, which were positioned between the poles of a 3470 GMW electromagnet (Fig. 3(c)). The measurements were performed at normal incidence. The reflected beam from the samples passed sequentially through a photoelastic modulator I/FS50 Hinds Instruments (PEM), a half-wave plate Thorlabs FR600HM (HWP), and a beam splitter Thorlabs WPA10 (BS), before being collected by a balanced photodetector Thorlabs PDB250A2 (BPD). In the MOKE setup, the combination of the photoelastic modulator with two lock-in amplifiers operating at 50 kHz (Signaloc 2100, Hinds Instruments) and 100 kHz (SR830 DSP, Stanford Research Systems), along with a balance detector, enabled the simultaneous measurement of Kerr rotation ($\theta$) and Kerr ellipticity ($\varepsilon$).
Data Availability
@@
Acknowledgment
S.S.J. and V.A. acknowledge the Academy of Finland (Project No. 356797). V.A. acknowledges the PREIN Flagship (Project No. 320167) and the Finnish Foundation for Technology Promotion. S. F. acknowledges a MURI project from the U. S. Army Research Office (Grant No. W911NF-24-1-0224). I.F., R.C., and A.D. acknowledge the Swedish Research Council for Sustainable Development (Formas) (Project No. 2021-01390), Thuréus Forskarhem och Naturminne Foundation, and the Swedish Research Council (VR) (Project No. 2024-05025). N.K. acknowledges financial support from the Magnus Ehrnrooth Foundation. We acknowledge the Micronova Nanofabrication Centre, supported by Aalto University, for providing resources for this work. We also thank Julius Hohlfeld and Junyoung Hyun for their assistance with the measurement setup.
@@ Acknowledgment
Competing interests
The authors declare no competing interests.

### v2

@@ II.3 Experimental realization of the optical Tellegen metasurface
The fabricated nanocones are generally symmetric, as can be seen in the top-view Fig. 3(a) (dark regions centered on the bright spots). Although some imperfections and slight tilts can be present in the cone geometry (Fig. 3(b)), they do not have preferred directions, and their statistical averaging would result in zero net chirality. Importantly, even if weak chirality were locally introduced due to such imperfections, it would not contribute to nonreciprocal optical rotation in our configuration. Indeed, a chiral layer placed on top of a mirror does not produce a net polarization rotation, as the rotation accumulated during forward propagation is compensated by the rotation accumulated upon reflection during backward propagation. Therefore, any residual chirality in the fabricated structures does not affect the observed Tellegen effect. In sharp contrast, the polarization rotation due to the Tellegen effect doubles after forward and backward propagations.
We measure the magneto-optical Kerr rotation angle ($\theta$) and Kerr ellipticity ($\epsilon$) using a magneto-optical setup illustrated in Fig. 3(c) (see more details in Methods and Supplementary Section 6). The samples were positioned between the poles of an electromagnet generating a magnetic field of $B=\pm 1.3$ T, perpendicular to the measured surface. This field alternately reversed the spontaneous magnetization of the nanocones along their axes. The obtained hysteresis curves can be found in Supplementary Section 7. The incident light was directed normally onto the metasurface, passing through a hole in the electromagnet. In the measurements of the Tellegen effect, magneto-optical data were acquired under an external magnetic field of $B=1.3$ T. This ensures that the meta-atoms in all three metasurfaces remain in the same uniformly saturated magnetization state, making the metasurfaces magneto-optically equivalent and rendering our extraction technique accurate. The results for the measurements of the Tellegen effect in the absence of an external field are available in Supplementary Section 7.
@@
Figures 3(d)–(f) present the experimental Kerr rotation and ellipticity for the three metasurfaces ($\theta_{1,2,3}$ and $\epsilon_{1,2,3}$) as functions of the free-space wavelength. For comparison, we also include the simulated Kerr parameters for periodic metasurfaces. It is important to note that these full-wave simulations cannot fully capture the complexity of the experimental metasurfaces, featuring meta-atoms with significant variations in their overall sizes (although with the same proportions) and quasi-random distances between neighboring meta-atoms. In the simulations, we assume a periodic arrangement of meta-atoms with a period of $a=330$ nm, using averaged geometric parameters of the meta-atoms depicted in the inset of Fig. 3b).
@@
Despite the inherent differences between the quasi-random experimental metasurfaces and their periodic simulated counterparts, we observe both qualitative and satisfactory quantitative agreement between the measured and simulated Kerr spectra. All the spectra exhibit resonant dispersion. For the first metasurface with a spacer thickness of $h_{1}$, the simulated Kerr parameters are slightly larger in amplitude than the measured ones, although the spectral correspondence remains strong. For the second metasurface with thickness $h_{2}$, the experimental and simulated data show good agreement both spectrally and quantitatively. However, for the third metasurface with $h_{3}$, the simulated results are redshifted by approximately 150 nm relative to the experimental data.
@@
By applying our proposed extraction technique (summarized in Fig. 2), we calculate the polarizabilities of the experimental metasurfaces using both measured and simulated data from Figs. 3(d)–(f). The extracted polarizability spectra are plotted in Figs. 3(g)–(i). One can observe a good qualitative agreement between the measured and simulated data, both spectrally and in terms of amplitude. All three effects—gyroelectric, gyromagnetic, and Tellegen—exhibit comparable amplitudes within the wavelength range of interest. Comparing these results with the simulated polarizabilities of the idealized meta-atom in Fig. 1, we observe similar response, although the resonance frequencies of the effects have shifted. These quantitative differences stem from the substantial geometric variations between the idealized meta-atom in Fig. 1(a) and the realistic meta-atom in Fig. 3(b). Furthermore, in our extraction, we model the array of meta-atoms as a homogenized sheet of electric and magnetic currents with deeply subwavelength thickness. This introduces an additional source of inaccuracy, as the actual meta-atom height is on the order of $\lambda/6$. Nevertheless, both the simulated and measured results provide strong evidence for the resonant nature of the diagonal Tellegen effect in the proposed optical metasurface. The magnitudes of the Tellegen polarizability $|\alpha_{\rm te}|$, extracted from both the measured data and the simulations for the idealized meta-atoms in Fig. 1 at 830 nm, reach approximately $6\times 10^{-33}$ ${\rm m}^{2}\cdot{\rm s}$. These results demonstrate that the observed amplitude of the Tellegen effect is comparable to that of the resonant gyroelectric effect present in the metasurface. This implies that the Tellegen effect is sufficiently strong to be relevant for magneto-optical applications. The measured Tellegen polarizability $|\alpha_{\rm te}|$ for the metasurface in the absence of an external field (in the remanent magnetization state) is around half of that for the saturated sample (see Supplementary Section 7). This reduction arises because the remanent magnetization does not reach complete saturation: the cobalt layer exhibits a polycrystalline and slightly irregular microstructure resulting from the deposition process, which limits the achievable uniformity of magnetic alignment. A higher remanent magnetization could potentially be achieved by improving the crystallinity of the cobalt segment, for instance, through a post-deposition magnetic annealing step. In the following section, we calculate and discuss the magnitude of the observed value and determine the corresponding bulk magnetoelectric parameter.
III Discussion
@@
To evaluate the macroscopic effective-medium parameters of the purely Tellegen colloid shown in Fig. 4(d), we apply the Maxwell-Garnett mixing rule [50, Sec. 7.2.2]. Here, we assume the host medium to be water, however, other dielectric media would yield comparable results, differing mainly in the effective permittivity of the colloid. The volumetric concentration of the meta-atoms is set to $N=3.47\times 10^{20}\penalty 10000\ {\rm m}^{-3}$, corresponding to an experimentally feasible volume fraction of the cones $N_{V}=0.1$. For our calculations, we use numerically extracted individual polarizabilities (which are independent of incident angle) of the idealized nanocone depicted in the inset of Fig. 1(a). It is worth noting that the dipolar approximation assumed in this work is fully adequate for determining the bulk Tellegen parameter $\chi_{\rm eff}$. This is because higher-order multipoles, even if present in individual meta-atoms, do not contribute to this parameter when the meta-atoms are randomly distributed in a dilute colloid [7].
Figure 4(e) presents the effective bulk calculated Tellegen parameter, $\chi_{\rm eff}$, exhibiting a Lorentzian-type resonant spectral shape. At resonance near 850 nm, $\chi_{\rm eff}$ reaches approximately $10^{-3}$, which is at least 100 times larger than values observed in natural materials such as chromium oxide [23].
We can further make a rough estimate of the magnitude of $\chi_{\rm eff}$ using experimental data from Fig. 3(i). Since the experimentally extracted collective Tellegen polarizability $\alpha_{\rm te}$ in Fig. 3(i) is of the same order as the simulated individual Tellegen polarizability [36], the resulting bulk Tellegen parameter $\chi_{\rm eff}$ would also be of the order of $10^{-3}$ ($\chi_{\rm eff}$ is linearly proportional to the Tellegen polarizability in the first-order approximation [50, Eq. (7.108)]).
Owing to its isotropic response, the Tellegen colloid exhibits identical magneto-optical behavior regardless of the illumination direction. Figure 4(f) plots these “isotropic” Kerr parameters for the colloid, showing qualitative similarity to those of the membrane in Fig. 4(c), as both systems exhibit a purely Tellegen response. The estimated Kerr parameters are also 2 orders of magnitude higher than those in natural Tellegen materials such as chromium oxide [23]. Indeed, the Kerr rotation and ellipticity depend linearly on the $\chi_{\rm eff}$ parameter [28, Eq. (3.143)], and therefore, have the same level of enhancement.
@@
In conclusion, we have demonstrated the first experimental realization of the optical Tellegen effect in an artificial metamaterial system, with the strength of around 100 times greater than that of natural materials. The Tellegen metasurface was produced using conventional materials and a well-established, highly-scalable nanofabrication, ensuring practicality for diverse applications. Since the presented optical metasurface also exhibits gyroelectric and gyromagnetic effects, we introduced an approach for independently extracting each of the three nonreciprocal effects. This technique is expected to be instrumental in the magneto-optical characterization of other nonreciprocal optical metasurfaces, particularly where symmetry permits multiple nonreciprocal effects. Experimental measurements, including those of magneto-optical Kerr parameters and nonreciprocal polarizabilities, demonstrate strong agreement with full-wave simulations, confirming the accuracy and reliability of the proposed method.
@@
The inherent scalability of the chosen bottom-up nanofabrication approach and strong remanent magnetization inside meta-atoms facilitate the straightforward extension of the proposed nonreciprocal optical Tellegen metasurface to other metamaterial systems, including bulk Tellegen metamaterials and colloids. We expect that with a minor modification of the fabrication process (introducing an additional polymer layer beneath the nanocones that can facilitate lift-off process for the nanocones into water solution), it would be possible to realize the Tellegen colloid. Such an approach would benefit from the scalability of hole-mask colloidal lithography, which enables the preparation of multiple cm-scale samples suitable for producing colloids with high meta-atom concentrations. The possible agglomeration of the magnetic nanocones inside the colloid can be avoided using a standard procedure applying static electric charge to the colloid to make the meta-atoms repel one another [61]. Furthermore, by tailoring the geometry and composition of meta-atoms, the resonant wavelength can be precisely tuned across the optical spectrum. Opening-up the ability to modulate the Tellegen metamaterial in space and/or time has profound implications for generating and analyzing effective axion fields— which are inherently nonuniform [29, 35, 16]— and other related phenomena [64, 42, 27, 58, 49].
IV Methods
IV.1 Simulation
@@ IV.2 Sample fabrication
IV.3 Magneto-optical characterization
The magneto-optical Kerr effect (MOKE) of the samples was characterized using an NKT SuperK EXW-12 supercontinuum laser connected to an acousto-optical tunable filter (AOTF), providing a wavelength range of 500–1400 nm. The AOTF employed in our setup has two outputs: VIS/NIR (500–900 nm) and NIR2 (800–1400 nm). A focusing doublet, Glan-Thompson polarizer, and silver D-shaped mirror were used to direct the light into the magnet poles and focus it onto the samples, which were positioned between the poles of a 3470 GMW electromagnet (Fig. 3(c)). The measurements were performed at normal incidence. The reflected beam from the samples passed sequentially through a photoelastic modulator I/FS50 Hinds Instruments (PEM), a half-wave plate Thorlabs FR600HM (HWP), and a beam splitter Thorlabs WPA10 (BS), before being collected by a balanced photodetector Thorlabs PDB250A2 (BPD). In the MOKE setup, the combination of the photoelastic modulator with two lock-in amplifiers operating at 50 kHz (Signaloc 2100, Hinds Instruments) and 100 kHz (SR830 DSP, Stanford Research Systems), along with a balance detector, enabled the simultaneous measurement of Kerr rotation ($\theta$) and Kerr ellipticity ($\epsilon$). To ensure reproducibility, we fabricated and characterized 3 samples for each Al2O3 thickness. Additionally, measurements were conducted at several regions on each sample. No significant differences in the optical response of the metasurfaces were observed, confirming good reproducibility of both the fabrication and measurement processes.
Acknowledgments
S.S.J. and V.A. acknowledge the Academy of Finland (Project No. 356797). S.S.J. acknowledges the Walter Ahlström Foundation and the Finnish Foundation for Technology Promotion. V.A. acknowledges the PREIN Flagship (Project No. 320167) and the Finnish Foundation for Technology Promotion. S. F. acknowledges a MURI project from the U. S. Army Research Office (Grant No. W911NF-24-1-0224). I.F., R.C., and A.D. acknowledge the Swedish Research Council for Sustainable Development (Formas) (Project No. 2021-01390), Thuréus Forskarhem och Naturminne Foundation, and the Swedish Research Council (VR) (Project No. 2024-05025). N.K. acknowledges financial support from the Magnus Ehrnrooth Foundation. We acknowledge the Micronova Nanofabrication Centre, supported by Aalto University, for providing resources for this work. We also thank Julius Hohlfeld and Junyoung Hyun for their assistance with the measurement setup.
Competing interests
The authors declare no competing interests.
Availability of data and materials